1. (2 Points) Explain clearly why the solution to the homogeneous system Ax = 0 with a nonsingular
coecient matrix is x = 0.

2. (2 Points) Under what conditions does a diagonal matrix D = diag(d1,d2,.....,dn) have an inverse
D^-1? What is the inverse D^-1 when these conditions are met? Justify your answers.

3. (3 Points) Let A be an mxn matrix and B be an nxm matrix where m > n. Show that the nxn matrix AB is not invertible.
1
Expert's answer
2013-02-18T11:52:48-0500
1. If A is nonsingular thendetA <>0 and there is A^(-1). So A^(-1)A=A^(-1)0 =>
x=A^(-1)0=0

2. If there is some F that DF=FD=I then F have to be too diagonal
and F = diag(f1,f2,.....,fn). Then we have that f_i * d_i =
1. So,
D is invertible iff each d_i is invertible, and then D^(-1)=diag{d1^(-1),
... , dn^(-1)}.

3. It is well known that if a is m by n matrix then rank(A) <=
min{m,n}. In our case we have that rank(A)<=n, rank(B)<=n.
also for any matrices we know that rank(AB)<=min{rank(A),rank(B)}
We have that AB will be m by m matrix and thus it will be invertible iff
det(AB)<>0 iff rank(AB)=m, but
rank(AB)<=min{rank(A),rank(B)}<=n<m so AB is not invertible.